### v2

# Universal properties of the evolution of the Universe in modified loop quantum cosmology

Abstract
In this paper, we systematically study the evolution of the Universe in the framework of a modified loop quantum cosmological model (mLQC-I) with various inflationary potentials, including chaotic, Starobinsky, generalized Starobinsky, polynomials of the first and second kinds, generalized T-models and natural inflation. In all these models, the big bang singularity is represented by a quantum bounce, and the evolution of the Universe both before and after the bounce is universal and weakly depends on the inflationary potentials, as long as the evolution is dominated by the kinetic energy of the inflaton at the bounce. In particular, the evolution in the pre-bounce region can be universally divided into three different phases: pre-bouncing, pre-transition, and pre-de Sitter. The pre-bouncing phase occurs immediately before the quantum bounce, during which the evolution of the Universe is dominated by the kinetic energy of the inflaton. Thus, the equation of state of the inflaton is about one, $w(\phi)\simeq 1$. Soon, the inflation potential takes over, so $w(\phi)$ rapidly falls from one to negative one. This pre-transition phase is very short and quickly turns into the pre-de Sitter phase, whereby the effective cosmological constant with a Planck size takes over and dominates the rest of the contracting phase. In the entire pre-bounce regime, the evolution of the expansion factor and the inflaton can be approximated by analytical solutions, which are universal and independent of the inflation potentials.
I Introduction
Since its incarnation in 1980 Guth (1981), the inflationary paradigm has achieved great success, resolving many long-standing problems of the standard big bang cosmology, and is consistent with all cosmological and astrophysical observations conducted so far Akrami et al. (2020). However, the paradigm is past incomplete Borde et al. (2003), Due to the Big Bang singularity, where all physical quantities become infinite. Hence, the initial conditions of inflation are usually imposed at a moment sufficiently far from the singularity but early enough so that the wavelengths of all the observational modes are within the Hubble horizon.
Then, a natural question arises: What happened before inflation? This important question is expected to be answered by quantum gravity (QG), a theory that has not yet been established, despite enormous efforts in the past century, although many candidates have been proposed. Among them are string/M-theory Green et al. (2012); Becker et al. (2006) and loop quantum gravity (LQG) Ashtekar and Lewandowski (2004); Thiemann (2007); Bojowald (2010); Gambini and Pullin (2011); Rovelli and Vidotto (2014a). In particular, in the last two decades, LQG has been rigorously applied to understand singularity resolution in various cosmological models (for recent reviews, see Refs. Ashtekar and Singh (2011); Elizaga Navascués and Marugán (2021); Li and Singh (2023); Agulló et al. (2023)). In all these models, a coherent picture of Planck scale physics has emerged: the big bang singularity is replaced by a quantum bounce, purely due to quantum geometric effects. This framework is often referred to as loop quantum cosmology (LQC).
The physical implications of LQC have also been studied using the effective descriptions of the quantum spacetime derived from coherent states Taveras (2008), whose validity has been numerically verified for various spacetime Singh (2018), especially for states that sharply peaked on classical trajectories at late times Kamiński et al. (2020). Effective dynamics provide a definitive answer to the resolution of the big bang singularity, replaced by a quantum bounce when the energy density of matter reaches a maximum value determined purely by the underlying quantum geometry Ashtekar and Singh (2011); Elizaga Navascués and Marugán (2021); Li and Singh (2023); Agulló et al. (2023).
Despite a wealth of results on the singularity resolution and phenomenology of the very early universe obtained in the framework of LQC, an important issue that has remained open is its connection to LQG (see, for example, Beetle et al. (2017); Bojowald (2021) for discussions). The starting point of LQC is to first classically reduce the Hamiltonian from infinitely many to a few gravitational degrees of freedom by imposing homogeneity and then to quantize the classically reduced Hamiltonian using the techniques of LQG. However, in LQG, the processes of symmetry reduction and quantization do not commute in general, and it is important to understand how well the physics of the full LQG is captured by LQC. In addition, in LQG, the Hamiltonian usually consists of two parts, the Euclidean and Lorentz parts, and one follows different processes to quantize each part Ashtekar and Lewandowski (2004); Thiemann (2007); Bojowald (2010); Gambini and Pullin (2011); Rovelli and Vidotto (2014a). However, in LQC, only the quantization of the Euclidean part was considered by taking advantage of the properties of the classical Hamiltonian, in which the Euclidean and Lorentz parts are proportional to each other for the flat Friedmann-Lemaítre-Robertson-Walker (FLRW) universe Ashtekar and Singh (2011).
In the past decade, the above issue has been extensively studied by both bottom-up and top-down approaches, from which an important conclusion is emerging: LQC and its major predictions are robust. In particular, the big bang singularity is resolved in the models studied so far. However, dramatic changes in the evolution of the universe in the pre-bounce phase are also found Li et al. (2021).
In the bottom-up approach Yang et al. (2009), symmetries are still imposed before quantization, but the Lorentzian term is treated independently by applying Thiemann’s regularization from the full theory of LQG Thiemann (1998a, b). In doing so, it was found that the resultant wave function is now described by a fourth-order difference equation Yang et al. (2009), which is referred to as mLQC-I Li et al. (2021). For a systematic derivation of the model, we refer readers to Assanioussi et al. (2018, 2019). For sharply peaked states, the resulting quantum dynamics are well described by effective Friedman-Raychaudhuri (FR) equations Li et al. (2018a), with which it was found that both the resolution of the big bang singularity Li et al. (2018b) and the existence of a subsequent desired slow-roll inflation are generic Li et al. (2019a).
In the top-down approach, using complexifier coherent states and treating the Euclidean and Lorentzian terms in the scalar constraint separately as in full LQG, Dapor and Liegener (DL) first obtained a modified LQC model Dapor and Liegener (2018a, b), which is quite similar to mLQC-I but with the $\mu_{0}$-scheme, as DL used a fixed graph Dapor and Liegener (2018a, b). To address this shortcoming, the DL work was extended to allow for graph-changing dynamics in an approach based on the path-integral reformulation of LQG Han and Liu (2021), whereby a consistent mLQC-I model can be obtained.
In this paper, we shall systematically study the evolution of the Universe in the framework of mLQC-I. Note that such studies have already been conducted in the post-bounce regime ($t\geq t_{B}$) in Li et al. (2018a, b, 2019a, 2021), where $t_{B}$ denotes the bounce time. It was found that the evolution in this regime is universal and independent of the inflationary potential, as long as the evolution is dominated at the quantum bounce by the kinetic energy of the inflaton
$\frac{1}{2}\dot{\phi}_{B}^{2}\gg V(\phi_{B}),$ (1.1)
where an over dot denotes the derivative with respect to the cosmic time, and $\phi_{B}$ is the value of the inflaton at the bounce. Initial conditions satisfying Eq.(1.1) are important because they always lead to slow-roll inflation Li et al. (2018a, b, 2019a, 2021), quick similar to that in LQC Bonga and Gupt (2016); Zhu et al. (2017). Therefore, in this paper, we focus on the pre-bounce regime ($t\leq t_{B}$).
Specifically, the rest of the paper is organized as follows: In Sec.II we briefly review the mLQC-I model and write down the dynamical Hamiltonian equations. To understand the major properties of mLQC-I, we also write down the corresponding FR equations. In Sec.III, we study the evolution of the Universe with several inflationary potentials, including chaotic, Starobinsky, generalized Starobinsky, polynomials of the first and second kinds, generalized T-models and natural inflation. In the last case, we consider several choices of parameters involved in the model. Note that the chaotic and natural inflationary models have some tensions with current observations Akrami et al. (2020). The reasons that we still consider them here are two-folds: First, after quantum gravitational effects are taken into account, such tensions can be alleviated Zhu et al. (2016). Second, the universal properties of the evolution of the universe both before and after the quantum bounce are independent of the choice of inflationary potentials in LQC Zhu et al. (2017); Shahalam et al. (2017, 2018); Sharma et al. (2018, 2019); Shahalam et al. (2020); Levy and O. Ramos (2024). We find that this is also true in mLQC-I. In all these models, we find that the evolution of the Universe in the pre-bounce regime ($t\leq t_{B}$) is universal and weakly depends on the inflationary potentials, as long as the kinetic-energy-dominated condition (1.1) is satisfied at the bounce. In all these models, the evolution is clearly divided into three different epochs, pre-bouncing, pre-transition, and pre-de Sitter, as defined by Eq(II), in terms of the equation of state of the inflaton
$w(\phi)\equiv\frac{P(\phi)}{\rho(\phi)},$ (1.2)
where $P(\phi)$ and $\rho(\phi)$ are the pressure and energy density of the inflationary field, respectively, given by Eq.(II). Moreover, in these epochs the expansion factor $a(t)$ and the inflaton $\phi(t)$ are universal and weakly depend on the initial conditions $\phi_{B}$ as well as on the inflationary potentials as shown explicitly by Fig. 31 of Sec. IV. In this section, we also show that $a(t)$ and $\phi(t)$ can be well-approximated by Eqs.(3.16) and (3.19), respectively, where $d_{n}$ and $e_{n}$ are constants determined by fitting the analytical expressions with their numerical solutions, as given in Sec. IV. Therefore, the evolution of the Universe in the pre-bounce regime ($t\leq t_{B}$) is also universal, quite similar to that in the post-bounce regime ($t\geq t_{B}$) Li et al. (2018a, b, 2019a, 2021). This is important and will significantly facilitate the studies of cosmological perturbations Li et al. (2020a, b, 2024), as already shown in LQC Zhu et al. (2017). The paper ends in Sec.V, in which our main results and concluding remarks are presented.
II Effective Dynamics of Modified Loop Quantum Cosmology
In this section, we provide a summary of the modified Friedmann dynamics for mLQC-I Li et al. (2021). Its dynamics can be obtained directly from the effective Hamiltonian, given by Yang et al. (2009)
$\mathcal{H}=\frac{3v}{8\pi G\lambda^{2}}\left\{\sin^{2}(\lambda b)-\frac{(\gamma^{2}+1)\sin^{2}(2\lambda b)}{4\gamma^{2}}\right\}+\mathcal{H_{M}},$ (2.1)
where $G$ is the Newtonian constant, $v\equiv v_{0}a^{3}$, and $v_{0}$ is the volume of a fiducial cell in the $\mathcal{R}^{3}$ spatial manifold, and $a$ is the expansion factor of the Universe. The variable $b$ denotes the momentum conjugate of $v$ and satisfies the canonical relation
$\{b,v\}=4\pi G\gamma,$ (2.2)
where $\gamma$ is known as the Barbero-Immirzi parameter whose value is set to $\gamma\approx 0.2375$ using black hole thermodynamics in LQG Meissner (2004). The parameter $\lambda$ is defined as $\lambda^{2}\equiv\Delta=4\sqrt{3}\pi\gamma\ell^{2}_{Pl}$, where $\Delta$ denotes the minimal area of the area operator in LQG Ashtekar and Lewandowski (2004); Thiemann (2007); Bojowald (2010); Gambini and Pullin (2011); Rovelli and Vidotto (2014b); Ashtekar and Pullin (2017).
In this paper, we consider only the case in which the matter is characterized by a single scalar field $\phi$ with potential $V(\phi)$, for which $\mathcal{H}_{M}$ is given by
${\mathcal{H}}_{M}=\frac{1}{2}v\left(\frac{p^{2}_{\phi}}{v^{2}}+2V(\phi)\right),$ (2.3)
where $p_{\phi}$ is the momentum of $\phi$. Then, the corresponding energy density and pressure of the scalar field are given by
$\rho$$\equiv$ $\frac{\mathcal{H_{M}}}{v}=\frac{1}{2}\dot{\phi}^{2}+V(\phi),$
$P$$\equiv$ $-\frac{\partial\mathcal{H}_{M}}{\partial v}=\frac{1}{2}\dot{\phi}^{2}-V(\phi).$ (2.4)
The basic variables $b$, $v$, $\phi$ and $p_{\phi}$ satisfy the Hamiltonian equations
$\dot{v}$ $=$ $\{v,\mathcal{H}\}$ (2.5)
$=$ $\frac{3v\sin{(2\lambda b)}}{2\gamma\lambda}\left\{(\gamma^{2}+1)\cos{(2\lambda b)}-\gamma^{2}\right\},$
$\dot{b}$ $=$ $\{b,\mathcal{H}\}=\frac{3\sin^{2}{(\lambda b)}}{2\gamma\lambda^{2}}\left\{\gamma^{2}\sin^{2}{(\lambda b)}-\cos^{2}{(\lambda b)}\right\}$ (2.6)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-4\pi G\gamma P,$
$\dot{\phi}$ $=$ $\{\phi,\mathcal{H}\}=\frac{p_{\phi}}{v},$ (2.7)
$\dot{p}_{\phi}$ $=$ $\{p_{\phi},\mathcal{H}\}=-vV_{,\phi},$ (2.8)
where $V_{,\phi}\equiv dV(\phi)/d\phi$. Eqs.(2.5) - (2.8) are the first-order ordinary differential equations for the four canonical variables ($v,b;\phi,p_{\phi}$). Once the initial conditions are specified at a given moment, for example, $t=t_{B}$, they uniquely determine the trajectory of the evolution of the Universe. Such initial conditions are often imposed at the quantum bounce Ashtekar and Singh (2011); Li et al. (2021), at which the expansion factor reaches its minimal value and the energy density reaches its maximum.
To see the above clearly, it is very suggestive to write the above four dynamical equations in the form of the modified FR equations Li et al. (2021)
$H^{2}$ $=$ $\frac{8\pi G\rho}{3}\left(1-\frac{\rho}{\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\right)\left(1+\frac{\gamma^{2}\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}{\left(\gamma^{2}+1\right)\left(1+\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\right)^{2}}\right),\;\;\;(t\geq t_{B}),$ (2.9)
$\frac{\ddot{a}}{a}$ $=$ $-\frac{4\pi G}{3}\left(\rho+3P\right)+\frac{4\pi G\rho^{2}}{3\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\left(\frac{\left(7\gamma^{2}+8\right)-4\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}\left(5\gamma^{2}+8\right)\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}}{\left(\gamma^{2}+1\right)\left(1+\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\right)^{2}}\right)$ (2.10)
$+4\pi GP\left(\frac{3\gamma^{2}+2+2\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}}{\left(\gamma^{2}+1\right)\left(1+\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\right)}\right)\frac{\rho}{\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}},\;\;\;(t\geq t_{B}),$
where $H\equiv\dot{v}/(3v)=\dot{a}/a$, and
$\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}\equiv\frac{\rho_{c}}{4(1+\gamma^{2})},\quad\rho_{c}\equiv\frac{3}{8\pi\lambda^{2}\gamma^{2}G}.$ (2.11)
It is evident from these two equations that the energy conservation law $\dot{\rho}+3H(\rho+P)=0$ holds. Substituting Eq.(II) into this equation, we find
$\quad\ddot{\phi}+3H\dot{\phi}+V_{,\phi}=0,$ (2.12)
while in terms of $\rho$ and $P$, we also have
$\dot{b}=-4\pi G\gamma(\rho+P)=-4\pi G\gamma\dot{\phi}^{2}.$ (2.13)
It should be noted that, Eqs.(2.9) and (2.10) hold only after the quantum bounce ($t\geq t_{B}$), as already indicated in these equations, at which we have $\rho(t_{B})=\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}$ and $H(t_{B})=0$, so the expansion factor reaches its minimal value $a_{B}\equiv a(t_{B})$. When $t\gg t_{B}$ (or equivalently, $\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}\ll 1$), Eqs.(2.9) and (2.10) reduce to their relativistic limits
$H^{2}$ $\simeq$ $\frac{8\pi G}{3}\rho,\;(t\gg t_{B}),$ (2.14)
$\frac{\ddot{a}}{a}$ $\simeq$ $-\frac{4\pi G}{3}\left(\rho+3P\right),\;(t\gg t_{B}).$ (2.15)
In particular, it is interesting to note that $\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}\simeq 10^{-12}$ at the onset of inflation Ashtekar and Singh (2011); Li et al. (2021). Therefore, during the inflationary phase, the modified FR equations are well approximated by its classical limits (2.14) and (2.15).
In the pre-bounce phase ($t\leq t_{B}$), the modified FR equations take the form Li et al. (2021)
$H^{2}$ $=$ $\frac{8\pi G_{\alpha}\rho_{\Lambda}}{3}\left(1-\frac{\rho}{\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\right)\left(1+\frac{\rho\left(1-2\gamma^{2}+\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\right)}{4\gamma^{2}\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}\left(1+\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\right)}\right),\;\;\;(t\leq t_{B}),$ (2.16)
$\frac{\ddot{a}}{a}$ $=$ $-\frac{4\pi G_{\alpha}}{3}\left(\rho+3P-2\rho_{\Lambda}\right)+4\pi G_{\alpha}P\left(\frac{2-3\gamma^{2}+2\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}}{\left(1-5\gamma^{2}\right)\left(1+\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\right)}\right)\frac{\rho}{\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}$ (2.17)
$-\frac{4\pi G_{\alpha}\rho^{2}\left(2\gamma^{2}+5\gamma^{2}\left(1+\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\right)-4\left(1+\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\right)^{2}\right)}{3\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}\left(1-5\gamma^{2}\right)\left(1+\sqrt{1-\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}}\right)^{2}},\;\;\;(t\leq t_{B}),$
where $G_{\alpha}\equiv\alpha G$, and
$\alpha\equiv\frac{1-5\gamma^{2}}{\gamma^{2}+1},\quad\rho_{\Lambda}\equiv\frac{3}{8\pi G\alpha\lambda^{2}(1+\gamma^{2})^{2}}.$ (2.18)
From Eqs.(2.16) and (2.17) we can see that at the bounce $\rho(t_{B})=\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}$, the universe contracts to its minimal volume $v=v_{0}a_{B}^{3}$ at $t=t_{B}$. Afterward, it smoothly passes to the expansion phase, but is now described by Eqs.(2.9) and (2.10). The smoothness is shown explicitly in Li et al. (2018a, b, 2019a), and can be also seen from Eqs.(2.12) and (2.13), which hold across the bounce.
When $t\ll t_{B}$ (or $\rho/\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}\ll 1$), Eqs.(2.16) and (2.17) reduce to
$H^{2}$ $\simeq$ $\frac{8\pi G_{\alpha}}{3}\rho_{\Lambda}\left(1-\frac{\rho}{\rho_{\Lambda}}\right),\;(t\ll t_{B}),$ (2.19)
$\frac{\ddot{a}}{a}$ $\simeq$ $\frac{8\pi G_{\alpha}}{3}\rho_{\Lambda}\left(1-\frac{\rho+3P}{2\rho_{\Lambda}}\right),\;(t\ll t_{B}),~{}~{}~{}~{}$ (2.20)
which are quite different from Eqs.(2.14) and (2.15). In particular, the effective Planck-scale cosmological constant $\rho_{\Lambda}$ soon dominates the evolution of the pre-bounce phase, whereby a de Sitter spacetime is obtained in the pre-bounce phase but with a Planck-scale cosmological constant $\rho_{\Lambda}\simeq{\cal{O}}(\rho_{\text{pl}})$. In addition, the Newtonian constant $G$ is replaced by $G_{\alpha}(=\alpha G)$, where $\alpha$ is defined by Eq.(2.18). More remarkably, this Planck-scale cosmological constant is filtered out by the quantum bounce and disappears miraculously after the bounce, whereby the classical FR equations are obtained, as shown explicitly by Eqs.(2.14) and (2.15). This is significantly different from LQC Ashtekar and Singh (2011), in which the evolution of the Universe is symmetric with respect to the bounce  [^2] More precisely, it is symmetric for kinetic energy-dominated initial conditions $\dot{\phi}_{B}^{2}\gg 2V(\phi_{B})$ Ashtekar and Singh (2011); Li et al. (2021)..
The background evolution of the universe in the post-bounce regime ($t\geq t_{B}$) has been studied extensively in Li et al. (2018a, b, 2019a, 2021), with various potentials, including chaotic, fractional monodromy, Starobinsky, and non-minimally coupling Higgs potentials. Among several remarkable features, it was found that the evolution of the universe is universal for any given potential, as long as the initial conditions imposed at the bounce are kinetic-energy dominated, that is
$\frac{1}{2}\dot{\phi}_{B}^{2}\gg V(\phi_{B}),$ (2.21)
where $\phi_{B}(\equiv\phi(t_{B}))$ is the value of the scalar field at the bounce. Moreover, evolution can be universally divided into three phases:
$(i)\;\text{bouncing}\;(w(t)\simeq 1),$
$(ii)\;\text{transition}\;(-1<w(t)<1),$
$(iii)\;\text{inflationary},\;(w(t)\simeq-1),$ (2.22)
as shown in Fig. 1. Here, the equation of state of the scalar field $w(t)$ is defined as follows:
$w(t)\equiv\frac{P(t)}{\rho(t)}=\frac{\frac{1}{2}\dot{\phi}^{2}-V}{\frac{1}{2}\dot{\phi}^{2}+V}.$ (2.23)
Figure 1: The evolution $w(t)$ of the equation of state in the post-bounce phase ($t\geq t_{B}$) for the chaotic potential with different initial conditions $\phi_{B}$ for $\dot{\phi}_{B}>0$. In this plot, $t_{B}$ is set to zero.
It should be noted that in Fig. 1, only the case of the chaotic potential was considered. However, the general behavior of the evolution of the universe is universal and independent of the choice of potential, as long as the initial conditions are kinetic energy dominated. These remarkable features are also shared by LQC, first found in Zhu et al. (2017) and later confirmed in various cases Shahalam et al. (2017, 2018); Sharma et al. (2018, 2019); Shahalam et al. (2020).
Moreover, during the bouncing phase, the expansion factor $a(t)$ and the scalar field $\phi(t)$ can be well described by the analytical solution Li et al. (2019a)
$a(t)$ $=$ $\left(1+24\pi\rho^{I}_{c}\left(1+\frac{A_{0}\gamma}{1+B_{0}t}\right)t^{2}\right)^{\frac{1}{6}},$
$\phi(t)$ $=$ $\phi_{B}+\text{sgn}\left(\dot{\phi}_{B}\right)\frac{\text{sinh}^{-1}{\sqrt{24\pi\rho^{I}_{c}\left(1+\frac{C_{0}\gamma^{2}}{1+D_{0}t}\right)t}}}{\sqrt{12\pi\left(1+\frac{C_{0}\gamma^{2}}{1+D_{0}t}\right)}},$
where $A_{0}=1.2$, $B_{0}=6$, $C_{0}=1.2$, and $D_{0}=2$ with relative errors no larger than $0.3\%$, As shown in Fig. 2. Again, the analytical solutions are universal and independent of the choice of inflationary potential $V(\phi)$ as long as the initial kinetic-energy dominated condition (1.1) holds.
Figure 2: The numerical and analytical solutions of $a\left(t\right)$ and $\phi(t)$ for different initial values of $\phi_{B}$ that result in at least 50 e-folds for the chaotic potential with $\dot{\phi}_{B}>0$. This universal behavior is true for any given inflation potential $V(\phi)$ as long as the condition (2.21) holds at the bounce.
In addition, it was shown that the probability of the occurrence of a desired slow-roll inflation is generic. In particular, it was found that for such a desired slow-roll inflation to not occur, the probability is Li et al. (2019a)
$P_{\;\text{mLQC-I}}(\text{not realized})\lesssim 1.12\times 10^{-5},$ (2.25)
where “desired” means the one that is consistent with current observations. This is comparable to $P_{\text{LQC}}(\text{not realized})\lesssim 2.74\times 10^{-6}$, obtained in LQC Ashtekar and Sloan (2011a, b).
Therefore, in this paper, we shall focus on the evolution of the universe in the pre-bounce regime ($t\leq t_{B}$) and pay particular attention to the universal properties of evolution.
III Numerical solutions of the Evolution of the Universe
Since the kinetic-energy dominated initial conditions imposed at the bounce always lead to slow-roll inflation for any given potential, in the rest of this paper, we shall consider only such initial conditions with several well-studied inflationary potentials Akrami et al. (2020); Kallosh and Linde (2022), including chaotic, Starobinsky, $\alpha$-attractive, and natural inflation potentials.
The advantage of imposing the initial conditions at the bounce is that $\dot{\phi}_{B}$ is determined uniquely up to a sign for any given initial data $\phi_{B}$ via the relation $\rho(t_{B})=\rho^{\text{I}}_{c}$, which yields
$\dot{\phi}_{B}=\pm\sqrt{2(\rho^{\scriptscriptstyle{\mathrm{I}}}_{c}-V(\phi_{B}))}.$ (3.1)
On the other hand, from Eqs.(2.9)-(2.10) and (2.16)-(2.17) we can see that these equations are scaling-invariant with respect to the expansion factor $a\rightarrow a/L_{o}$. Therefore, without loss of generality, we can always set $a_{B}=1$. Then, the initial conditions are reduced to the choice of
$\left(\phi_{B},\text{sgn}\left(\dot{\phi}_{B}\right)\right).$ (3.2)
In addition, using the translation invariance $t\rightarrow t+t_{0}$, in the rest of this paper, we shall set $t_{B}=0$.
To process further, following Li et al. (2019b, 2018b, a, 2021), let us first introduce the following physical quantities:
  • The first-order Hubble rate and potential slow-roll parameters
    $\epsilon_{H}$ $=-\frac{\dot{H}}{H^{2}},$ $\eta_{H}=-\frac{\ddot{H}}{2H\dot{H}},$ (3.3a)
    $\epsilon_{V}$ $=\frac{1}{16\pi G}\left(\frac{V_{,\phi}}{V}\right)^{2},$ $\eta_{V}=\frac{V_{,\phi\phi}}{8\pi GV}.$ (3.3b)
    These sets of slow-roll parameters are typically used for different purposes Baumann (2011). In particular, the slow-roll parameters with the subscript “$V$” can be used to determine which part of the potential can successfully drive inflation. On the other hand, slow-roll parameters with subscript “$H$” are used for numerical simulations to define when slow-roll inflation begins and ends. In the classical regime, the scale factor acceleration equation satisfies the following relation:
    $\ddot{a}=aH^{2}(1-\epsilon_{H}).$ (3.4)
    The Universe experiences an accelerated expansion when $\epsilon_{H}<1$, whereas slow-roll inflation occurs only when $\epsilon_{H}\ll 1$ and $\eta_{H}\ll 1$ Baumann (2011). For the sake of concreteness, we define the onset of inflation as the time $t_{i}$ when $|\eta_{H}|=0.03$ for the first time in the transition phase. The end of the slow-roll inflation is defined at the time $t_{\text{end}}$ when $|\epsilon_{H}|=1$ for the first time after $t_{i}$.
  • The e-fold $N$ during the inflationary phase: This number is usually defined as follows:
    $N=\ln\left({\frac{a(t_{\text{end}})}{a(t_{i})}}\right).$ (3.5)
    To have a successful slow-roll inflation, the inflation potential has to be very flat, so that the Universe can expand large enough Baumann (2011). All the cosmological problems can be resolved if the Universe expands about 60 e-fold during the inflationary phase, although its exact value depends on the inflationary models.
With the above in mind, we are now ready to solve the four dynamical equations (2.5) - (2.8) for any given initial conditions (3.2)  [^3] It is found that numerically it is more convenient to solve the dynamical equations (2.5) - (2.8) than the ones of Eqs.(2.9)-(2.10) and (2.16)-(2.17), although they should give the same results. However, in the latter, the integration needs to be carried out in the pre- and post-bounce separately, and then connect them smoothly across the bounce.. In the following, we shall study the chaotic, Staribinsky, $\alpha$-attractor and natural inflation potentials, separately.
III.1 Chaotic Potential
The chaotic inflationary potential is given by Linde (1983)
$V(\phi)=\frac{1}{2}m^{2}\phi^{2},$ (3.6)
where the mass is set to $m=1.26\times 10^{-6}\;m_{pl}$ Akrami et al. (2020). Since it has already been shown that the occurrence of a desired slow-roll inflation is generic Li et al. (2019a), it is sufficient for us to find the initial conditions that lead to $N\geq 45$ e-folds.
In this model, the modified Friedmann and Klein-Gordon equations are invariant under the replacement
$(\phi,\dot{\phi})\rightarrow(-\phi,-\dot{\phi}),$ (3.7)
rendering it sufficient to consider only half of the parameter space with $\dot{\phi}_{B}>0$. Here, $\phi_{B}$ can assume any value in the range $|\phi_{B}|\leq\phi^{I}_{\text{max}}=\sqrt{2\rho^{I}_{c}}/m$, where $\phi^{I}_{\text{max}}$ is determined by $V(\phi^{I}_{\text{max}})=\rho^{I}_{c}$ Li et al. (2019a).
For any chosen initial condition $\phi_{B}$ with $\dot{\phi}_{B}>0$, which yields approximately 60 e-folds [cf. Fig. 3], we numerically solve the dynamical equations (2.5) - (2.8), and find that the evolution of the universe in the post-bounce regime ($t\geq 0$) can always be divided into three different phases, bouncing, transition and inflation, as shown in Fig. 1. In Fig. 2, three numerical solutions are shown together with the analytical solution given by Eq.(II) during the bouncing phase, $t\in(0,10^{4})\;t_{\text{pl}}$. Note that the universal behavior of the evolution of the universe during the bouncing phase does not depend on the specific choice of $\phi_{B}$, as long as the kinetic-energy dominated condition (1.1) is satisfied at the bounce Li et al. (2019a).
Figure 3: The e-folds $N$ during the inflationary period for the chaotic potential (3.6) with $m=1.26\times 10^{-6}\;m_{pl}$ for $\dot{\phi}_{B}>0$. The case $\dot{\phi}_{B}<0$ can be obtained by the symmetry $(\phi,\dot{\phi})\rightarrow(-\phi,-\dot{\phi})$, a particular property owned by this model.
With the same initial conditions imposed at the bounce, we numerically solve the dynamic equations (2.5)- (2.8) for $t\leq 0$ and obtain the solutions in the pre-bounce regime. In Fig. 4 (a), the corresponding equation of state $w(t)$ defined by Eq.(2.23) is plotted. From this figure, we can see that the evolution of the universe in the pre-bounce phase can be also universally divided into three different phases:
$(i)\;{\text{p}re-bouncing}\;(w(t)\simeq 1),$
$(ii)\;{\text{p}re-transition}\;(-1<w(t)<1),$
$(iii)\;{\text{p}re-de\;Sitter}\;(w(t)\simeq-1).$ (3.8)
Figure 4: (a) The evolution of the equation of state in the pre-bounce phase for the chaotic potential with various choices of $\phi_{B}$ for $\dot{\phi}_{B}>0$. (b) Ratio between the kinetic energy and potential of the scalar field. (c) The ratio $\rho(t)/\rho_{\Lambda}$ between the energy density of the scalar field and the effective cosmological constant, together with the Hubble parameter $H(t)$.
The transition occurs at $t\simeq-10t_{\text{pl}}$ and is very rapid. From Fig. 4 (b), we can see that this is because the kinetic energy of the scalar field decreases dramatically as time becomes more and more negative and is soon dominated by its potential energy. In addition, the effective cosmological constant quickly dominates the evolution of the universe $\rho(t)/\rho_{\Lambda}\ll 1$, as shown in Fig. 4 (c), whereby the Hubble parameter becomes a (negative) constant, and spacetime quickly approaches de Sitter. As a result, the pre-bouncing phase is also very short compared with the pre-de Sitter phase.
Figure 5: The numerical solutions of $a(t)$ and $\phi(t)$ for the chaotic potential with $\dot{\phi}_{B}>0$. In the figures, the numerical curves with different initial conditions are indistinguishable.
In Fig. 5, the numerical solutions of $a(t)$ and $\phi(t)$ are plotted for three different initial conditions, which all satisfy (1.1). Form this figure it can be seen that the numerical solutions are indistinguishable. In fact, they are not only indistinguishable with different initial conditions, but also indistinguishable with different potentials, as long as the condition (1.1) holds at the bounce, as to be shown below.
III.2 Starobinsky Potential
The Starobinsky potential results from adding an $R^{2}$ term to the Einstein-Hilbert action Starobinskii (1979). After a conformal transformation, in the Einstein frame, the theory is equivalent to a scalar field with the potential Maeda (1988)
$V(\phi)=\frac{3m^{2}}{32\pi G}\left(1-\exp\left(-\sqrt{\frac{16\pi G}{3}}\phi\right)\right)^{2},$ (3.9)
where the mass is set to $2.49\times 10^{6}m_{\text{pl}}$ Akrami et al. (2020). Unlike the case with chaotic potential, now the slow-roll inflation can take place only when $\phi>\phi_{\text{end}}\simeq 0.188m_{\text{pl}}$ Li et al. (2019a).
Refer to caption Refer to caption
Figure 6: E-folds produced during slow inflation for different values of $\phi_{B}$ for the Starobinsky potential (3.9). Left-panel: $\dot{\phi}_{B}>0$. Right-panel: $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: The numerical solutions $a(t)$ and $\phi(t)$ with indicated initial conditions for the Starobinsky potential (3.9) in the post-bounce regime ($t>0$), where (a) For $\dot{\phi}_{B}>0$, and (b) for $\dot{\phi}_{B}<0$. To compare with the analytical solutions given by Eq.(II), we also plot them out, denoted by the red solid lines.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 8: The equation of state for the Starobinsky potential (3.9) for different values of $\phi_{B}$ that result in greater than 50 e-Folds for (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: The numerical solutions of $a(t)$ and $\phi(t)$ with various initial values of $\phi_{B}$ for the Starobinsky potential (3.9), where (a) for $\dot{\phi}_{B}>0$ and (b) For $\dot{\phi}_{B}<0$.
In addition, the Starobinsky potential does not share the symmetry of the chaotic one, given by Eq.(3.7). Therefore, in the following, we consider both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$ separately. In particular, in Fig. 6 we plot the e-fold $N$ of inflation defined by Eq.(3.5) vs different initial conditions $\phi_{B}$ for $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively.
To confirm the universality of the evolution of the universe during the post-bounce regime ($t\geq 0$), in Fig. 7 we plot out $a(t)$ and $\phi(t)$ with the initial conditions
[…]
so that during slow-roll inflation, about 60 e-folds can be produced.
On the other hand, with the same initial conditions as those chosen in Fig. 7, we plot the equation of state $w(t)$ in the post-bounce ($t\geq 0$) and the pre-bounce ($t\leq 0$) regimes in Fig. 8. From these figures, it can be clearly seen that in the post-bounce regime, three different phases defined by Eq.(II) exist. The only difference from the chaotic potential case is that the transition phase now occurs at approximately $t\simeq 10^{5}\;t_{\text{pl}}$, instead of $t\simeq 10^{4}\;t_{\text{pl}}$ as shown in Fig. 1.
Equally remarkable, the evolution of the universe in the pre-bounce regime ($t\leq 0$) is also universally divided into three phases defined by Eq.(III.1). However, the time $t_{m}$ defined by $\dot{H}(t_{m})\simeq 0$ in the pre-transition phase weakly depends on the initial values of $\phi_{B}$ but can be significantly different with different signs of $\dot{\phi}_{B}$, as shown in Fig. 8, where $t_{m}\simeq-5t_{\text{pl}}$ for $\dot{\phi}_{B}>0$, and $t_{m}\simeq-10.5t_{\text{pl}}$ for $\dot{\phi}_{B}<0$.
In Figs. 9, the numerical solutions of $a(t),\;\phi(t)$ for several different choices of initial conditions $\phi_{B}$ with $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$ are plotted out, respectively. From the figures we can see that the scalar field $\phi(t)$ quickly approaches a constant, which is negative for $\dot{\phi}_{B}>0$ and positive for $\dot{\phi}_{B}<0$.
III.3 $\alpha$- attractor Potentials
Refer to caption Refer to caption
Figure 10: E-folds for different values of $\phi_{B}$ for the generalized Starobinsky potentials (3.11) with $\alpha=0.0962$ and $V_{0}=10^{-12}$ GeV.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 11: The equation of state for the generalized Starobinsky potentials (3.11) for different values of $\phi_{B}$ that result in at least 50 e-Folds with $\alpha=0.0962$ and $V_{0}=10^{-12}$ GeV. (a) for $\dot{\phi}_{B}>0$ and (b) for $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 12: The numerical solutions of $a(t)$ and $\phi(t)$ for the $\alpha$-attractor potential (3.11) with $\alpha=0.0962$ and $V_{0}=10^{-12}$ GeV. (a) For $\dot{\phi}_{B}>0$, and (b) for $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Figure 13: E-fold $N$ for different values of $\phi_{B}$ for the polynomial potential of the first kind given by Eq.(3.12) where $V_{0}=3.3787*10^{-15}$ GeV and $\mu=0.31075$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 14: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the polynomial potential of the first kind given by Eq.(3.12) with $V_{0}=3.3787\times 10^{-15}$ GeV and $\mu=0.31075$. (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 15: The numerical solutions of $a(t)$ and $\phi(t)$ with different initial values of $\phi_{B}$ for the polynomial potential of the first kind given by Eq.(3.12), where $V_{0}=3.3787\times 10^{-15}$ GeV and $\mu=0.31075$. (a) For $\dot{\phi}_{B}>0$ and (b) for $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Figure 16: E-fold $N$ for different values of $\phi_{B}$ for the polynomial potential of the second kind given by Eq.(3.13) with $V_{0}=3.2599\times 10^{-15}$ GeV and $\mu=0.01043$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 17: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the polynomial potential of the second kind given by Eq.(3.13) with $V_{0}=3.2599\times 10^{-15}$ GeV and $\mu=0.01043$, where (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 18: The numerical and analytical solutions of $a(t)$ and $\phi(t)$ with different initial values of $\phi_{B}$ for $\dot{\phi}_{B}>0$ and the polynomial potential of the second kind given by Eq.(3.13), where $V_{0}=3.2599\times 10^{-15}$ GeV and $\mu=0.01043$.
Refer to caption Refer to caption
Figure 19: E-folds $N$ for the generalized T-Models given by (3.14) with $V_{0}=1.96\times 10^{-15}$ and $\alpha=0.0962$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 20: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the the generalized T-Models given by (3.14) with $V_{0}=1.96\times 10^{-15}$ and $\alpha=0.0962$, where (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 21: The numerical solutions of $a(t)$ and $\phi(t)$ for various initial conditions of $\phi_{B}$ for the generalized T-Models given by (3.14) with $V_{0}=1.96*10^{-15}$ and $\alpha=0.0962$. (a) For with $\dot{\phi}_{B}>0$ and (b) for $\dot{\phi}_{B}<0$.
The $\alpha$-attractor potentials usually take several different forms Akrami et al. (2020). In this subsection, we consider several representative cases.
III.3.1 Generalized Starobinsky Potentials
One kind of the $\alpha$-attractor potentials from supergravity takes the form Ferrara et al. (2013)
$V(\phi)=V_{0}\left(1-e^{-\frac{\phi}{\sqrt{6}\alpha}}\right)^{2},$ (3.11)
which is often referred to as the generalized Starobinsky potentials or E-models Kallosh and Linde (2022). For $\alpha=\sqrt{9/(8\pi G)}$, the above potentials are precisely reduced to the Starobinsky potential given by Eq.(3.9).
In the following, we consider the case $\alpha=0.0962$ with $V_{0}=10^{-12}$ GeV. Then, we find that the initial values of $\phi_{B}$ for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, with which the e-fold $N$ is always no less than 45, as shown in Fig.10. With several such values of $\phi_{B}$, in Fig. 11 the corresponding equation of state $w(t)$ for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$ are plotted out. From these figures, we can see that each of the post-bounce and pre-bounce regimes can be universally divided into three phases, as given in Eqs.(II) and (III.1).
In Fig. 12, we plot the numerical solutions of $a(t)$ and $\phi(t)$ for several initial values of $\phi_{B}$ for $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively. From these figures, it can be seen that, similar to the previous cases, the numerical solutions are almost indistinguishable from each other.
III.3.2 Polynomial of the first kind
The polynomial potential of the first kind takes the form Kallosh and Linde (2022)
$V(\phi)=V_{0}\frac{\phi^{4}}{\phi^{4}+\mu^{4}},$ (3.12)
where $V_{0}$ and $\mu$ are two constants. In this paper, we take the values of $V_{0}$ and $\mu$ as $V_{0}=3.3787\times 10^{-15}$ GeV and $\mu=0.31075$ Bhattacharya et al. (2023). Then, in Figs. 13 and 14 we plot the e-fold $N$ and the equation of state $w(t)$ for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$.
Similar to the previous cases, the evolution of the universe in each of the pre- and post-bounce regimes can be universally divided into three epochs, describing by Eqs.(II) and (III.1), respectively.
In Fig. 15, we plot the functions $a(t)$ and $\phi(t)$ with various initial values of $\phi_{B}$ for $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively.
III.3.3 Polynomial of the second kind
The polynomial potential of the second kind is given by Kallosh and Linde (2022)
$V(\phi)=V_{0}\frac{\sqrt{\phi^{2}+\mu^{2}}-\mu}{\sqrt{\phi^{2}+\mu^{2}}+\mu}.$ (3.13)
In this paper, we adopt the best fitting values of $V_{0}$ and $\mu$ with current observations given by $V_{0}=3.2599\times 10^{-15}$ and $\mu=0.01043$ Bhattacharya et al. (2023).
In Figs. 16 and 17 we plot $N(\phi_{B})$ and the corresponding equation of state $w(t)$ for several initial conditions $\phi_{B}$ for $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively. These values of $\phi_{B}$ guarantee that the e-folds during inflation will be about 60 as shown in Fig. 16. On the other hand, Fig. 17 shows again that the evolution of the universe in each of the pre- and post-bounce regimes can be universally divided into three different epochs.
In Fig. 18 we plot numerical solutions of $a(t)$ and $\phi(t)$. Again, these numerical solutions are almost indistinguishable with different initial conditions as specified in the figures, which show again the universality of the evolution of the Universe.
III.3.4 Generalized T-models
The generalized T-model potentials of the superconformal inflationary models are given by Kallosh et al. (2013)
$V(\phi)=V_{0}\tanh^{2}\left({\frac{\phi}{\sqrt{6}\alpha}}\right),$ (3.14)
where the best fitting values of $V_{0}$ and $\alpha$ with current cosmological observations are given by $V_{0}=1.96\times 10^{-15}$ and $\alpha=0.0962$ Bhattacharya et al. (2023).
In Figs. 19 and 20 we plot $N(\phi_{B})$ and the corresponding equation of state $w(t)$ for several initial conditions $\phi_{B}$ with $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively. These values of $\phi_{B}$ also guarantee that the e-folds of the expansions of the universe during the inflationary period will be about 60, as shown in Fig. 20. On the other hand, Fig. 20 shows again that the evolution of the universe in each of the pre- and post-bounce regimes can be universally divided into three different epochs, as described by Eq.(II) in the post-bounce regime and by Eq.(III.1) in the pre-bounce regime.
In Fig. 21, we plot the numerical solutions of $a(t)$ and $\phi(t)$ with several initial conditions for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$ case, which all satisfy the condition (1.1) at the bounce. Again, they show the universality of the solutions.
III.4 Natural Inflation
The Natural inflationary potential, motivated from a pseudo-Nambu-Goldstone boson, takes the form Adams et al. (1993),
$V\left(\phi\right)=\Lambda^{4}\left[1+\cos{\left(\frac{\phi}{\mu}\right)}\right],$ (3.15)
where $\mu\equiv f/{\cal{N}}$, $\Lambda\simeq m_{\text{GUT}}\simeq 10^{16}$ GeV, $f\simeq m_{\text{pl}}$, and ${\cal{N}}$ is a dimensionless constant. In the following, we consider only three representative cases, $\mu/m_{\text{pl}}=\left(0.3,\;1,\;5\right)$.
III.4.1 $\mu=0.3\;m_{\text{pl}}$
In this case, we first identify the values of $\phi_{B}$ with both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, for which the e-folds can be approximately 60, as shown in Fig. 22.
Refer to caption Refer to caption
Figure 22: E-folds of natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=0.3m_{\text{pl}}$ for different values of the initial conditions $\phi_{B}$, where (a) for the case $\dot{\phi}_{B}>0$ and (b) for the case $\dot{\phi}_{B}<0$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 23: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=0.3m_{\text{pl}}$, where (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 24: The numerical solutions of $a(t)$ and $\phi(t)$ for various initial conditions of $\phi_{B}$ for the natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=0.3$, where (a) for $\dot{\phi}_{B}>0$, and (b) for $\dot{\phi}_{B}>0$.
On the other hand, in Fig. 23 the equation of state is plotted with some of these initial values (with which the corresponding e-folds are about 60). From the figure, it can be seen again that the evolution of the universe is universal in both the pre- and post-bounce regimes, and in each of which three distinguishable epochs exist, as described by Eqs.(II) and (III.1).
The corresponding numerical solutions of $a(t)$ and $\phi(t)$ are shown in Fig. 24 for $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively, form which it cna be seen clearly the universality of the numerical solutions.
Refer to caption Refer to caption
Figure 25: E-folds of natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=m_{\text{pl}}$ for different values of the initial conditions $\phi_{B}$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 26: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=1\;m_{\text{pl}}$, where (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 27: The numerical solutions of $a(t)$ and $\phi(t)$ for various initial conditions of $\phi_{B}$ for the natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=1m_{\text{pl}}$, where (a) for $\dot{\phi}_{B}>0$ and (a) for $\dot{\phi}_{B}<0$.
III.4.2 $\mu=m_{\text{pl}}$
In this subsection, similar to the previous ones, we first identify the initial conditions $\phi_{B}$ for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, with which the e-folds of inflation are approximately 60. In Fig. 25 we show two sets of such data. Note that the number of such sets is infinitely large because the potential is now a periodic function of $\phi$. With some of the initial conditions presented in Fig. 25, we plot the corresponding equation of state $w(t)$ in Fig. 26, which again shows the universal properties described by Eqs.(II) and (III.1).
Conversely, the corresponding numerical solutions of $a(t)$ and $\phi(t)$ are given in Fig. 27 for $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively.
Refer to caption Refer to caption
Figure 28: E-folds of natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=5m_{\text{pl}}$ for different values of the initial conditions $\phi_{B}$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 29: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=5\;m_{\text{pl}}$, where (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 30: The numerical solutions of $a(t)$ and $\phi(t)$ for various initial conditions of $\phi_{B}$ with $\dot{\phi}_{B}>0$ for natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=5m_{\text{pl}}$.
III.4.3 $\mu=5\;m_{\text{pl}}$
In this case, in Fig. 28 we plot the e-folds of the natural inflation for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, whereby we can read off the initial conditions of $\phi_{B}$ for which the inflation will last about 60 e-folds.
Then, with several choices of such initial values of $\phi_{B}$ presented in Fig. 28, the equation of state $w(t)$ is plotted in Fig. 29. Again, we can clearly identify the three distinguishable epochs in each of the pre- and post-bounce regimes.
The corresponding numerical solutions of $a(t)$ and $\phi(t)$ are plotted in Fig. 30 for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 31: The numerical solutions of $a(t)$ and $\phi(t)$ for various potentials as specified in the figures. To see clearly the difference among the different models, we group the plot into two groups: Left Panels: Models with chaotic, Starobinsky and natural inflationary potentials. Right Panels: Models with the polynomials of the first and second kinds and the generalized T-models, although they are almost indistinguishable among all these models. (a) The evolution of the expansion factor $a(t)$ for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$. (b) The evolution of the scalar field $\phi(t)$ for $\dot{\phi}_{B}>0$. (c) The evolution of the scalar field $\phi(t)$ for $\dot{\phi}_{B}<0$.
IV Analytical Solutions for the evolution of the Universe
As shown previously Li et al. (2019a), the evolution of the Universe in the post-bounce regime ($t\geq t_{B}$) is universal, and can be divided into three different epochs defined as in Eq.(II). In addition, during the bouncing phase, the solutions of $a(t)$ and $\phi(t)$ are given by the analytical solutions of Eq.(II) [Also see Fig. 2].
In this section, we are going to show that this is also the case in the pre-bounce regime ($t\leq t_{B}$). To show this, let us first plot the numerical solutions of $a(t)$ and $\phi(t)$ together for all the cases considered in the last section. Since for each potential the numerical solutions are indistinguishable for different initial values of $\phi_{B}$, we plot only one solution for each model. In addition, to see clearly the difference among the models, we also plot them in two groups, one for chaotic, Starobinsky and natural inflaitonary potentials and the other for the polynomials of the first and second kinds and the generalized T-models, although they are indistinguishable among all these models, as can be seen from Fig. 31, denoted by the right-hand and left-hand panels, respectively. The numerical solutions of $a(t)$ is independent of the signs of $\dot{\phi}_{B}$ and are given by Fig. 31 (a), while The numerical solutions of $\phi(t)$ depend on the signs of $\dot{\phi}_{B}$ and are given respectively by Figs. 31 (b) and (c). From Fig. 31 (a) we can see that the universe soon enters the pre-de Sitter phase, and $a(t)$ increases exponentially $a(t)\propto e^{-H_{\Lambda}t}$ as $t$ becomes more and more negative, where $H_{\Lambda}\equiv\sqrt{{8\pi\alpha G\rho_{\Lambda}/3}}$ [cf. Eq.(2.19)]. On the other hand, the evolution of the scalar field indeed depends on the signs of $\dot{\phi}_{B}$, as in the post-bounce regime, as cna be seen from Eq.(II). In particular, it quickly approaches a negative constant for $\dot{\phi}_{B}>0$ and a positive constant for $\dot{\phi}_{B}<0$, as shown explicitly by Figs. 31 (b) and (c).
To analytically model the expansion factor $a(t)$ in the pre-bounce regime, we find that it is convenient to assume that it takes the form
$a(t)=\begin{cases}\left(1+d_{0}\rho^{I}_{c}t^{2}\right)^{\frac{1}{6}}\sum_{n=2}^{4}{d_{n}t^{n}},&t_{m}\leq t\leq t_{B},\cr a(t_{m})\exp\left\{-H_{\Lambda}\left(t-t_{m}\right)\right\},&t\leq t_{m},\cr\end{cases}$ (3.16)
where $t_{m}$ is determined by $\dot{H}(t_{m})\simeq 0$, and $d_{n}\;(n=0,...,4)$ are fitting constants, which must satisfy the junction conditions at both the bounce ($t=0$) and the turning point ($t=t_{m}$). In particular, we require that $a(t)$ and its first derivative be contiguous across these two points, where $a(t)$ is given by Eq.(II) in the post-bounce regime ($t\geq 0$). As shown by Fig. 31, the curve of $a(t)$ weakly depends on the initial values of $\phi_{B}$ and the inflationary potentials, so we can fit Eq.(3.16) with any given inflationary potential and initial values of $\phi_{B}$, as long as the condition (1.1) is satisfied. In particular, for the Starobinsky potential with $\dot{\phi}_{B}>0$ we find
$d_{0}$ $\simeq$ $74.057,\;\;\;d_{2}\simeq 0.104,$
$d_{3}$ $\simeq$ $0.001,\;\;\;d_{4}\simeq 0.002,$ (3.17)
for which the relative errors $\delta a(t)$ are less than $1\%$ at any given moment $t\leq t_{B}$, as shown in Fig. 32, where the relative errors are defined by
$\delta{A}\equiv\left|\frac{A_{n}-A_{a}}{A_{n}}\right|,$ (3.18)
where $A_{n}$ and $A_{a}$ denote the numerical and analytical values of $A$ with $A=(a,\;\phi)$.
In addition, to model the scalar field $\phi(t)$, we find that in the whole pre-bounce regime ($t\leq t_{B}$), the scalar field can be cast in the form
$\phi(t)=e_{3}+\text{sgn}\left(\dot{\phi}_{B}\right)(e_{0}+e_{2}t)e^{e_{1}t},\;(t\leq t_{B}),$ (3.19)
where $e_{m}\;(m=0,...,3)$ are the fitting constants. Again, we require that $\phi(t)$ and its first derivative be continuous across the bounce, where $\phi(t)$ is given by Eq.(II) in the post-bounce regime ($t\geq t_{B}$). Then, we find that with the choices of fitting constants as
$e_{0}$ $\simeq$ $0.353,\;\;\;e_{1}\simeq 0.981,$
$e_{2}$ $\simeq$ $0.943,\;\;\;e_{3}\simeq-1.704,$ (3.20)
for $\dot{\phi}_{B}>0$, the relative errors $\delta\phi(t)$ are less than $1\%$ at any given moment $t\leq t_{B}$, as shown in Fig. 32. On the other hand, for $\dot{\phi}_{B}<0$, the coefficients $e_{n}$ are given by
$e_{0}$ $\simeq$ $0.353,\;\;\;e_{1}\simeq 0.981,$
$e_{2}$ $\simeq$ $0.094,\;\;\;e_{3}\simeq 3.802$ (3.21)
It should be noted that
$e_{3}=\phi_{\infty}-\phi_{B},$ (3.22)
$e_{0}=\text{sgn}\left(\dot{\phi}_{B}\right)\left(2\phi_{B}-\phi_{\infty}\right),$
where $\phi_{\infty}\equiv\phi(t=-\infty)$. Therefore, $e_{0}$ and $e_{3}$ depends on the choice of the initial values of $\phi_{B}$ as well as the sign of $\dot{\phi}_{B}$, as can be seen from Eqs.(IV) and (IV). However, once the sign of $\dot{\phi}_{B}$ is chosen, the coefficient $e_{3}$ is independent of the inflationary potentials, as can be seen from Figs. 31 and 32.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 32: The numerical and analytical solutions of $a(t)$ and $\phi(t)$ with various initial values of $\phi_{B}$ for the Starobinsky potential (3.9), where (a) for $\dot{\phi}_{B}>0$ and (b) For $\dot{\phi}_{B}<0$. The analytical solutions are given by Eqs.(3.16) - (IV).
V Conclusions and Remarks
In this paper, we have systematically studied the evolution of the Universe before the quantum bounce ($t\leq t_{B}$) in one of the most promising modified theories of loop quantum cosmologies, the so-called mLQC-I Li et al. (2021), which can be obtained by both bottom-up Yang et al. (2009); Assanioussi et al. (2018, 2019) and top-down Dapor and Liegener (2018a, b); Han and Liu (2021) approaches, as a complementary to the studies in the post-bounce regime ($t\geq t_{B}$) carried out recently in Li et al. (2018a, b, 2019a), where $t_{B}$ denotes the moment that the quantum bounce occurs, replacing the classical Big bang singularity. In this paper, we have mainly focused on the initial conditions for which the kinetic energy of the inflationary field $\phi$ dominates the evolution of the Universe at the bounce $t=t_{B}$, as given by Eq.(1.1). This set of initial conditions is important and leads generically to slow-roll inflation, as shown explicitly in Li et al. (2018a, b, 2019a). Such considerations do not lose their generality, as slow-roll inflation is a generic feature in this model, as shown in Li et al. (2019a).
With the above in mind, we have studied several important inflationary potentials tested recently by the Planck 2018 data Akrami et al. (2020). These include chaotic, Starobinsky, generalized Starobinsky, polynomials of the first and second kinds, generalized T-models, and natural inflation. Despite the fact that the chaotic and natural inflationary models have some tensions with current observations Akrami et al. (2020), in this paper, we found that the evolution of the Universe weakly depends on the specific form of the potentials, as long as the initially kinetic-energy dominated condition (1.1) holds at the quantum bounce.
In particular, we found that the evolution before the bounce can be universally divided into three different epochs, the pre-bouncing, pre-transition, and pre-de Sitter epochs, as defined in Eq.(III.1). Although the exact moment of the transition point defined by $\dot{H}(t_{m})=0$ depends on the form of the inflationary potentials, the division of the above three different epochs is sharp and clear, as shown by the equation of state $w(\phi)$ in each of the cases considered in Sec. III. In addition, the effective cosmological constant $\rho_{\Lambda}$ defined in Eq.(2.18) has Planck size and soon dominates the evolution of the Universe as one departs from the bounce to the deep contracting phase ($t\rightarrow-\infty$). Therefore, the pre-bouncing and pre-transition epochs are all very short compared with the pre-de Sitter epoch.
Moreover, the numerical solutions with the initial conditions that satisfy Eq.(1.1) considered in all the models of this paper are universal, as shown explicitly by Fig. 31, and can be approximated well by analytical solutions cast in the forms (3.16) for the expansion factor $a(t)$ and (3.19) for the inflation $\phi(t)$, where the constants $d_{n}$ and $e_{n}$ appearing in these expressions cna be obtained by fititng the analytical solutions with the numerical ones. For the expansion factor $a(t)$ we have found that the fitting ocnstants $d_{n}$ are given by Eq.(IV) for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$. However, for the scalar field, the fitting constant $e_{n}$ depend on the signs of $\dot{\phi}_{B}$, and are given respectively by Eqs.(IV) and (IV). This is because these coefficients depend on the choice of the initial values of $\phi_{B}$ and the signs of $\dot{\phi}_{B}$, as shown by Eq.(3.22).
The above remarkable features are also shared in the evolution of the Universe in the post-bounce regime not only in the framework of mLQC-I Li et al. (2018a, b, 2019a, 2021) but also in that of LQC Zhu et al. (2017); Shahalam et al. (2017, 2018); Sharma et al. (2018, 2019); Shahalam et al. (2020). We believe that this universal behavior of background evolution will significantly simplify the studies of cosmological perturbations in mLQC-I, as has already been the case in LQC Zhu et al. (2017).
Acknowledgements.
J.S. and C.B. are supported by the Baylor Physics graduate program, and R.P. and A.W. are partially supported by the US NSF grant: PHY-2308845.